# LDI: Localized Data Imputation

### v1

@@ VI-G Limited-Resource Setting
As shown in Table IX, LDI consistently outperforms the baseline across all datasets. These results show that LDI’s improvements are not dependent on large-scale language models; rather, its design effectively enhances performance through better input selection by providing the model with cleaner, more focused input. When smaller LLMs such as Llama 3.2 3B are used, focusing on localized input becomes especially important as these models have limited capacity to process and reason over lengthy inputs. If irrelevant attributes or uninformative examples are included, they can dilute the clues and distract the model from important patterns.
@@
TABLE IX: Performance comparison when deployed with a smaller language model (Llama 3.2 3B)
@@
VII Conclusion and Future Work
@@
Acknowledgments
This research was supported by Natural Sciences and Engineering Research Council of Canada and Alberta Innovates Graduate Student Scholarship program.
AI-Generated Content Acknowledgment
AI writing assistant was used for polishing text and checking the grammar.

### v2

@@ 6.9. Limited-Resource Setting
As shown in Table 9, LDI consistently outperforms the baseline across all datasets. These results show that LDI’s improvements are not dependent on large-scale language models; rather, its design effectively enhances performance through better input selection by providing the model with cleaner, more focused input. When smaller LLMs such as Llama 3.2 3B are used, focusing on localized input becomes especially important as these models have limited capacity to process and reason over lengthy inputs. If irrelevant attributes or uninformative examples are included, they can dilute the clues and distract the model from important patterns.
@@
Table 9. Performance comparison when deployed with a smaller language model (Llama 3.2 3B)
@@
7. Limitations
While our approach effectively captures meaningful dependencies for imputation, it also has some limitations. First, the current implementation supports textual and categorical attributes and treats numerical values as textual tokens. This design choice was made because detecting patterns embedded in textual fields was the main challenge, requiring specialized handling of linguistic and symbolic variations. As a result, purely numerical relationships (e.g., arithmetic dependencies) are not explicitly modeled. Second, the method discovers dependencies through repetition patterns among tuples. This behavior is not unique to our framework, as all FD-based methods inherently depend on repeated evidence to infer dependencies. Nonetheless, we further introduced a group sampling strategy to relax this requirement, allowing the discovery of dependencies in cases where repetition is limited.
8. Conclusion and Future Work
@@
Acknowledgements.
This research is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC). Additionally, Soroush Omidvartehrani is supported by the Alberta Innovates Graduate Student Scholarship (AI GSS).
Artifacts
The artifacts required to reproduce our experiments are publicly available at:
The repository provides the source code, datasets, and detailed instruction for experiment reproduction.